Grade 12Mechanics

a cylinder of mass M and radius 2R has a groove of radius R. a block of mass m is hung from the circumference of the cylinder and the cylinder itself is hung from a rigid support by a thread wound on the groove . if the acceleration of the centre of mass of the cylinder is a1, acceleration of the block is a2 downward and angular acceleration of th cylinder is b,the the relation between a1, a2 and b is

dolly nambi
9 Years agoGrade 12

1 Answer

Rituraj Tiwari
5 Years ago

To understand the relationship between the accelerations of the cylinder and the block, we need to analyze the motion of both bodies and how they interact with each other through the groove and the thread. Let’s break this down step by step.

Analyzing the System

We have a cylinder with mass M and radius 2R, which has a groove of radius R. A block of mass m is hung from the circumference of the cylinder, and the cylinder is suspended from a rigid support by a thread wound around the groove. Understanding the dynamics of this system involves looking at the forces and the resulting accelerations.

Defining Accelerations

Let’s denote:

  • a1: the acceleration of the center of mass of the cylinder.
  • a2: the downward acceleration of the block.
  • b: the angular acceleration of the cylinder.

Relationship Between Linear and Angular Accelerations

The key point here is that the motion of the block and the cylinder are interconnected. When the block descends with an acceleration a2, it unwinds the thread from the groove, causing the cylinder to rotate. The relationship between the linear acceleration of the cylinder and its angular acceleration is given by:

a1 = b * r

In this case, since the radius of the groove is R, we can express the relationship as:

a1 = b * R

Applying Newton's Second Law

Next, we apply Newton's second law to the block and the cylinder:

  • For the block: The forces acting on the block are its weight (mg) and the tension (T) in the thread. Thus, we can write:
  • m * a2 = mg - T

  • For the cylinder: The tension T creates a torque (τ) around its axis of rotation. The torque can be expressed as:
  • τ = T * R

    This torque is also related to the angular acceleration of the cylinder:

    τ = I * b

    Where I is the moment of inertia of the cylinder. For a solid cylinder, I = (1/2) * M * (2R)² = 2MR².

Combining Equations

From the torque equation, we have:

T * R = (2MR²) * b

Now, substituting T from the first equation:

m * a2 = mg - (2MR² * b / R)

Simplifying this gives us:

m * a2 = mg - 2M * R * b

Now, we can express b in terms of a1 and a2 using the relationship established earlier. We know that:

b = a1 / R

Substituting this into the equation yields:

m * a2 = mg - 2M * R * (a1 / R)

Thus, we arrive at:

m * a2 = mg - 2M * a1

Final Relationship Between Accelerations

Rearranging this equation gives us a direct relationship:

2M * a1 + m * a2 = mg

From here, we can derive various expressions depending on the specific values of M, m, and g. This relationship crucially links the downward acceleration of the block to the motion of the cylinder, demonstrating the interplay between linear and angular dynamics in this system.

In summary, the accelerations of the block and cylinder are deeply interconnected through the forces and torques acting on them, leading to the relationship we derived. This understanding is essential in problems involving rotational dynamics and linear motion.